OpenStudy (anonymous):

What is the approximate circumference of a circle with diameter 14 centimeters? Use 22/7 to approximate . answer: 11 cm 44 cm 22 cm 33cm

OpenStudy (anonymous):

c=pi*d c=22/7*14 c=44

OpenStudy (anonymous):

Hi. Thanks!

OpenStudy (anonymous):

np